OpenStudy (j3nny0rtiz):

please help... A 60 kg zebra is traveling 9 km/hr east. What is the zebra’s momentum? kg-m/s

OpenStudy (steve816):

Momentum = mass • velocity

OpenStudy (steve816):

Now, do you think you can figure it out?

OpenStudy (j3nny0rtiz):

540

OpenStudy (sshayer):

1 km=1000 m 1 hr =3600 s \[km/hr=\frac{ 1000 }{ 3600 }\] \[v=9 km/hr=\frac{ 9 \times 1000 }{ 3600 }=\frac{ 5 }{ 2 }~m/s\] momentum=mass*velocity \[=60 \times \frac{ 5 }{ 2 }=? kg~ m/s\]

OpenStudy (steve816):

Ooops, I totally forgot about the units. Listen to @sshayer , smart guy.

OpenStudy (j3nny0rtiz):

150 ?

OpenStudy (j3nny0rtiz):

@sshayer